# Generalizing Graph Transformers Across Diverse Graphs and Tasks via Pre-Training on Industrial-Scale Data

### v3

@@ 5.3. Friend Recall Task
Table 5. Linear probing results on Friend Recall Task. No-pretrain represents a random-initialized model without any self-supervised pre-training.
@@
Table 6. Results of fine-tuning the pre-trained model on Friend Recall Task. No-pretrain represents a random-initialized model without any self-supervised pretraining.
@@
Task description. The friend recall task is designed to encourage active players to invite their friends who have churned, thereby rewarding both parties with certain incentives. An active player in the game will see a list of their recently inactive friends. They can choose to click and invite them to play together in exchange for specified rewards. When the invitation is sent, the invited player’s social media account will receive a message. If the invited player chooses to accept the invitation, they will also receive corresponding rewards. Essentially, the Friend Recall Task is about ranking the churned players among their friends to maximize the probability of successful recall. With the friendships between players, we can model it as a link prediction task on graphs.
Setup. We build our datasets using friend recall records from Game C and Game D across a span of 5 consecutive months. The graph for each month can be regarded as a snapshot of friend recall records for a particular game. If an active player X invites a churned player friend Y, and player Y accepts the invitation, then the edge between them is labeled as a positive edge. Conversely, if the churned player Y does not accept the invitation, or if the active player X did not even invite the churned player friend Y, then the edge between them is labeled as negative. We train on the first three snapshots, validate on the fourth, and test on the final snapshot. Table 4 provides a summary of the datasets. The number of edges and the number of nodes are cumulative across the five snapshots.
We select all the self-supervised pre-trained models mentioned earlier as baseline methods. For linear probing setting, we still adopt ClusterGCN (Chiang et al., 2019) as the sampling algorithm to scale baseline methods to large graphs in inference. For the ease of mini-batch training on node pairs in the link prediction task, we choose Neighbor Sampling (Hamilton et al., 2017) as the sampling algorithm to scale baseline methods to large graphs under fine-tuning setting. We also compared some methods that do not utilize graph structures, such as XGBoost (Chen and Guestrin, 2016), MLP, and Bilinear. Similar to the node classification task, we conduct experiments under both the linear probing and fine-tuning settings for self-supervised methods. Following the majority of link prediction methods, we randomly sample node pairs within a batch as negative samples to train this classifier. We use the commonly used ranking metrics Hits@k and Mean Reciprocal Rank (MRR). We run the experiments 10 times with random seeds.
Results. Table 5 shows the results of our method and baseline methods under linear probing setting. Firstly, on both datasets, PGT achieves better results than all self-supervised learning baselines. This further underscores that our designed pre-training strategy can bring benefits to various downstream tasks at both the node and edge levels. Secondly, it can be observed that our method, along with most baseline methods, brings performance improvements over randomly initialized backbone networks through pre-training.
Table 6 reports the results under fine-tuning setting. We can observe that our proposed PGT framework consistently outperforms the baseline methods or achieves comparable results on both datasets. However, such improvements are not as significant as under linear probing setting. One possible reason is that we adopt Neighbor Sampling to support pairwise mini-batch training for GAT-based algorithms instead of ClusterGCN. However, the limitation of Neighbor Sampling is that it leads to training inefficiencies when extended to web-scale graphs (Zeng et al., 2019; Chiang et al., 2019). These results partially support our argument about the importance of designing consistent and efficient sampling algorithms for both the training and inference phases on web-scale graphs in industrial scenarios.
5.4. Ablation Studies
We further conduct a series of ablation studies to verify the contributions of the designs in PGT under the linear probing setting.
Table 7. Ablation studies of PGT pre-training objectives. We report ROC-AUC for Game A and B, MRR for Game C and D.
@@ 5.4. Ablation Studies
Table 8. Ablation studies of key designs in inference. We report ROC-AUC for Game A and B, MRR for Game C and D.
@@
Effect of pre-training objectives. Table 7 shows the results, where the ”w/o feature recon.” represents that we only employ local graph structure reconstruction objective in pre-training. Conversely, ”w/o local graph recon.” signifies that only the feature reconstruction objective is used. It can be observed that for two downstream tasks, both proposed pre-training objectives contribute to performance. Comparatively, the removal of the feature reconstruction objective results in a larger performance degradation, implying the significance of node features for graph learning tasks.
Effect of key designs in inference. Table 8 shows the influence of key proposed designs in inference. It can be observed that our proposed method of employing PPR sampling for inference attains performance comparable to that achieved using all neighbor nodes. This substantiates the efficacy of PPR in selecting informative nodes without the necessity for a large receptive field. The notation ”w/o proposed feature aug.” represents the absence of our proposed strategy of reusing the decoder for feature augmentation. Functioning as a plug-and-play component, it effectively improves performance across all datasets.
Efficiency analysis. Fig 2 compares inference performance and time of different sampling methods, using the same number of nodes to be predicted (test set on each dataset) and the same hardware conditions. It can be observed that our method delivers the most favorable outcomes in both performance and speed. It is noteworthy that, in contrast to the efficiency of ClusterGCN sampling during the training phase, it does not exhibit advantages in inference. A possible explanation could be that the nodes requiring predictions constitute only a small fraction of all nodes. The GNN still needs to perform inference across all sub-graphs to obtain all prediction results. This further substantiates the significance of decoupling the sampling of labelled and auxiliary nodes.
@@
Figure 2. Efficiency analysis of different sampling strategies in inference. ”Ours” refers to our graph transformer that uses PPR sampling, while the remaining four refer to GAT using conventional sampling strategies.
6. Performance on Public Benchmarks
PGT is designed as a general graph pre-training framework. We further conduct experiments on public benchmarks to substantiate its broad applicability.
6.1. Pre-training
Table 9. Statistics of public datasets.
Datasets #Nodes #Edges #Features
Cora 2,708 5,429 768
PubMed 19,717 44,338 768
ogbn-arxiv 169,343 1,166,243 768
ogbn-papers100M 111,059,956 1,615,685,872 768
Dataset. We conduct pre-training on a large paper citation network ogbn-papers100M (Hu et al., 2021). In this dataset, each node represents a paper and each directed edge indicates a citation. Each node is associated with its natural language title and abstracts. We leverage DeBERTa (He et al., 2020) to generate a 768-dimensional feature for each node. The statistics are listed in Table 9.
Baselines. The baselines are consistent with those on Tencent’s data in Section 3. To maximize the performance of the baselines, we adopt widely used GraphSAINT (Zeng et al., 2019) as the subgraph sampling method in our baselines. For all baseline methods, we employ GAT (Veličković et al., 2017) as the backbone network.
6.2. Node Classification Task
Table 10. Linear probing results on Node Classification Task. We report accuracy(%) for all datasets. No-pretrain represents a random-initialized model without any self-supervised pretraining.
Methods Cora PubMed ogbn-arxiv
MLP 30.76$\pm$0.76 53.50$\pm$0.55 42.43$\pm$1.23
No-pretrainGAT 48.21$\pm$0.53 62.56$\pm$0.88 67.12$\pm$0.27
DGI 32.44$\pm$0.64 54.34$\pm$0.76 46.24$\pm$0.34
GRACE 38.25$\pm$0.53 56.76$\pm$0.89 50.24$\pm$0.41
BGRL 45.43$\pm$0.34 61.99$\pm$1.02 68.08$\pm$0.33
GraphMAE 50.21$\pm$0.56 61.99$\pm$1.01 68.12$\pm$0.29
GraphMAE2 50.89$\pm$0.53 61.81$\pm$0.91 68.24$\pm$0.41
No-pretrainTransformer 48.67$\pm$0.43 62.63$\pm$1.22 67.45$\pm$0.43
PGT 52.43$\pm$0.46 62.87$\pm$1.89 70.22$\pm$0.42
Setup. In this study, we adopt Cora (McCallum et al., 2000), PubMed (Sen et al., 2008) and ogbn-arxiv (Hu et al., 2021), three popular citation network benchmarks for node classification. In these datasets, each node represents a paper and each directed edge indicates a citation. Each node is associated with its natural language title and abstracts. In alignment with the pre-training graph, we leverage DeBERTa (He et al., 2020) to generate a 768-dimensional feature for each node. The statistics are listed in Table 9. For Cora and PubMed, we follow the public data splits as (Veličković et al., 2018; Hou et al., 2022; He et al., 2023). For ogbn-arxiv, we follow the official splits in (Hu et al., 2021). We select all the self-supervised pre-trained models mentioned earlier as baseline methods. To maximize the performance of the baselines, we employ full neighborhood aggregation in inference. We evaluate our method under linear probing settings.
Results. Table 10 shows the results of our method and baselines under linear probing setting. We can observe that our proposed PGT framework generally outperforms all baselines in all three datasets. This substantiates that PGT is a general and effective graph pre-training framework, capable of applicability across a diverse range of graph domains. We also observe that some baseline methods manifest severe cases of negative transfer. We hypothesize that a potential reason could be that the citation networks themselves do not possess numerical features in the same feature space. The process of encoding textual features using pre-trained language models may introduce unavoidable noise. The noise in node features adversely impacts the generalization capability of the pre-trained models. Our proposed method for feature augmentation, which utilizes reconstructed features based on masked autoencoders, can be regarded as a form of feature denoising (Vincent et al., 2008). During inference, the model applies learned error-correction ability to the unmasked input, benefiting from the robustness learned during the training phase (Vincent et al., 2010).
7. conclusion

### v4

@@ VI-A Pre-training
TABLE IX: Linear probing results on Friend Recall Task. No-pretrain represents a random-initialized model without any pre-training.
@@
Improv. vs Best Baseline (%) 4.80% 2.68% 3.18% 4.20% 2.88% 3.48% 0.55% 1.01% 0.23% 1.77%
TABLE X: Results of fine-tuning the pre-trained model on Friend Recall Task. No-pretrain represents a random-initialized model without any pre-training.
@@
Improv. vs Best Baseline (%) 0.76% 0.78% 0.36% 0.20% 2.10% 0.49% -0.92% -0.75% 0.24% 0.14%
Dataset. We clarify some fundamental concepts in the online gaming business. Tencent’s online games and social media platforms are closely interconnected. Players can log into various games using their social platform accounts, such as WeChat, and can add new game friends within the games. For our pre-training dataset, we select game players on the WeChat platform as nodes and the union of each player’s friend relationships in a few games as edges. For privacy reasons, we extract a relatively large subgraph from the raw data. All subsequently mentioned datasets are constructed in this manner. We construct a 146-dimensional numerical feature for each node. These features include information about the account’s login history, activity duration, and in-game item purchase records over a certain period. Detailed information is shown in Table V.
Baselines. We compare our PGT framework with state-of-the-art graph pre-training and self-supervised learning algorithms, DGI [velivckovic2018deep], GRACE [zhu2020deep], BGRL [thakoor2021bootstrapped], GraphMAE [hou2022graphmae] and GraphMAE2 [hou2023graphmae2]. Other methods are not included in the comparison because they are not scalable to large graphs, e.g., GMAE [zhang2022graph]. Given the enormous scale of the pre-training graph, sampling algorithms like GraphSAINT [zeng2019graphsaint] or Neighbor Sampling [hamilton2017inductive] require more than 72 hours for pre-training. We adopt ClusterGCN [chiang2019cluster] as the subgraph sampling method in our baselines due to its superior training efficiency. For all baseline methods, we employ GAT [velivckovic2017graph] as the backbone.
VI-B Minor Detection Task
Task description. In accordance with relevant national regulations, Tencent’s online games have restrictions on the playing time of minors. The objective of the minors detection task is to identify users in a specific online game who appear to be minors. In this paper, we model it as a binary classification task per node on the graph. For each node to be predicted, we classify it as either an adult or a minor.
Setup. We built our datasets from the accounts of Game A and B, as illustrated in Table V. For both datasets, the training, validation, and test sets consist of 100,000, 100,000, and 800,000 nodes, respectively.
We select all the self-supervised pre-trained models mentioned earlier as baseline methods. In alignment with pre-training, we employ ClusterGCN as the subgraph sampling algorithm to scale baseline methods to large graphs. Following most graph self-supervised learning algorithms [hou2023graphmae2], we evaluate our method under two settings: linear probing and fine-tuning. For the linear probing setting, we freeze the pre-trained encoder and infer node representations. Then, using the node representations as input, we train a linear classifier to obtain the predictions. The purpose of the linear probing setting is to evaluate the pre-trained model’s ability to generate good node representations. For the fine-tuning setting, we attach a linear classifier to the last layer of the pre-trained encoder and then conduct supervised end-to-end fine-tuning. The purpose of the fine-tuning setting is to evaluate the pre-trained model’s ability to transfer its learned knowledge to unseen graphs and downstream tasks in low-label scenarios. For both settings, we run the experiments 10 times with random seeds.
Results. Table VI shows the linear probing results. We observe the following: (1) PGT significantly outperforms all graph self-supervised learning baselines, demonstrating its ability to generate more discriminative representations on unseen graphs. (2) Some pre-trained models exhibit negative transfer, performing worse than the non-pretrained GAT, especially on Game B. This suggests that knowledge learned during pre-training may not effectively transfer to unseen graphs. (3) Comparing backbones, our graph transformer with PPR sampling significantly outperforms GAT with ClusterGCN sampling. This is likely because ClusterGCN loses over 70-90% of edges when partitioning large graphs, causing a severe loss of structural information.
Table VII reports the fine-tuning results. In low-label scenarios, all pre-trained models show improvements compared to the non-pretrained model, demonstrating that pre-training provides better weight initialization for resource-scarce conditions. However, with sufficient labels, some pre-training methods fail to bring consistent improvements, possibly due to overfitting to noise in the training data, which reduces generalization. Across most experimental conditions, PGT consistently outperforms all baselines, demonstrating its adaptability to the widely used pretrain-finetune paradigm.
TABLE XI: Ablation studies of PGT pre-training objectives. We report ROC-AUC for Game A and B, MRR for Game C and D.
@@ VI-B Minor Detection Task
TABLE XII: Ablation studies of key designs in inference. We report ROC-AUC for Game A and B, MRR for Game C and D.
@@
VI-C Friend Recall Task
Task description. The friend recall task is designed to encourage active players to invite their friends who have churned, thereby rewarding both parties with certain incentives. An active player in the game will see a list of their recently inactive friends. They can choose to click and invite them to play together in exchange for specified rewards. When the invitation is sent, the invited player’s social media account will receive a message. If the invited player chooses to accept the invitation, they will also receive corresponding rewards. Essentially, the Friend Recall Task is about ranking churned players among their friends to maximize the probability of successful recall. With the friendships between players, we can model it as a link prediction task on graphs.
Setup. We build our datasets using friend recall records from Game C and Game D across a span of 5 consecutive months. The graph for each month can be regarded as a snapshot of friend recall records for a particular game. If an active player X invites a churned player friend Y, and player Y accepts the invitation, then the edge between them is labeled as a positive edge. Conversely, if the churned player Y does not accept the invitation, or if the active player X did not even invite the churned player friend Y, then the edge between them is labeled as negative. We train on the first three snapshots, validate on the fourth, and test on the final snapshot. Table VIII provides a summary of the datasets. The number of edges and the number of nodes are cumulative across the five snapshots.
We select all the self-supervised pre-trained models mentioned earlier as baseline methods. For the linear probing setting, we still adopt ClusterGCN [chiang2019cluster] as the sampling algorithm to scale baseline methods to large graphs in inference. For the ease of mini-batch training on node pairs in the link prediction task, we choose Neighbor Sampling [hamilton2017inductive] as the sampling algorithm to scale baseline methods to large graphs under fine-tuning setting. We also compared some methods that do not utilize graph structures, such as XGBoost [chen2016xgboost], MLP, and Bilinear. Similar to the node classification task, we conduct experiments under both the linear probing and fine-tuning settings for self-supervised methods. Following the majority of link prediction methods, we randomly sample node pairs within a batch as negative samples to train this classifier. We use the commonly used ranking metrics Hits@k and Mean Reciprocal Rank (MRR). We run the experiments 10 times with random seeds.
Results. Table IX shows the results of our method and baseline methods under linear probing setting. Firstly, on both datasets, PGT achieves better results than all self-supervised learning baselines. This further underscores that our designed pre-training strategy can bring benefits to various downstream tasks at both the node and edge levels. Secondly, it can be observed that our method, along with most baseline methods, brings performance improvements over randomly initialized backbone networks through pre-training.
Table X reports the results under fine-tuning setting. We can observe that our proposed PGT framework consistently outperforms the baseline methods or achieves comparable results on both datasets. However, such improvements are not as significant as under linear probing setting. One possible reason is that we adopt Neighbor Sampling to support pairwise mini-batch training for GAT-based algorithms instead of ClusterGCN. However, the limitation of Neighbor Sampling is that it leads to training inefficiencies when extended to web-scale graphs [zeng2019graphsaint, chiang2019cluster]. These results partially support our argument about the importance of designing consistent and efficient sampling algorithms for both the training and inference phases on web-scale graphs in industrial scenarios.
@@ VI-C Friend Recall Task
Figure 6: Efficiency analysis of different sampling strategies in inference. “Ours” refers to our graph transformer that uses PPR sampling, while the remaining four refer to GAT using conventional sampling strategies.
TABLE XIII: Comparison of different feature augmentation strategies. We report node classification accuracy (%).
Method Cora PubMed ogbn-arxiv
PGT w/ proposed feature aug. 70.11 69.24 70.22
PGT w/o augmentation 68.94 68.45 69.85
PGT w/ feature dropout (10%) 68.35 67.98 69.12
PGT w/ Gaussian noise 67.92 67.53 68.77
TABLE XIV: Comparison of different feature augmentation strategies on public datasets. We report node classification accuracy (%).
Feature Strategy Cora PubMed ogbn-arxiv
[…]
TABLE XV: PPR computation time for different datasets.
Dataset Number of Seed Nodes PPR Computation Time Infrastructure
Pre-training Graph 54M (10% of all nodes) $\sim$8 hours Spark+Angel (200 workers)
Game A 1M (labeled nodes) $\sim$30 minutes Spark+Angel (200 workers)
Game B 1M (labeled nodes) $\sim$25 minutes Spark+Angel (200 workers)
Game C 0.7M (labeled nodes) $\sim$20 minutes Spark+Angel (200 workers)
Game D 0.5M (labeled nodes) $\sim$15 minutes Spark+Angel (200 workers)
ogbn-papers100M 11M (10% of all nodes) $\sim$2 hours Single machine (128 cores)
ogbn-arxiv 0.17M (all nodes) $\sim$10 minutes Single machine (128 cores)
TABLE XVI: Sensitivity analysis of mixing coefficient $\lambda$.
$\lambda$ Value Cora PubMed ogbn-arxiv
0.01 $69.78\pm 0.52$ $68.92\pm 1.32$ $70.01\pm 0.57$
0.05 70.14 $\pm$ 0.48 $69.09\pm 1.27$ $69.87\pm 0.45$
0.1 $70.11\pm 0.44$ 69.24 $\pm$ 1.21 70.22 $\pm$ 0.42
0.2 $70.06\pm 0.51$ $69.17\pm 1.25$ $70.13\pm 0.49$
0.5 $70.10\pm 0.57$ $68.89\pm 1.31$ $69.95\pm 0.54$
TABLE XVII: Sensitivity analysis of mask rate.
Mask Rate Cora PubMed ogbn-arxiv
0.65 69.32 $\pm$ 0.63 68.57 $\pm$ 1.33 69.42 $\pm$ 0.58
0.75 69.78 $\pm$ 0.51 68.89 $\pm$ 1.29 69.85 $\pm$ 0.49
0.85 70.11 $\pm$ 0.44 69.24 $\pm$ 1.21 70.22 $\pm$ 0.42
0.90 69.87 $\pm$ 0.49 69.01 $\pm$ 1.27 69.96 $\pm$ 0.47
0.95 69.15 $\pm$ 0.61 68.43 $\pm$ 1.36 69.27 $\pm$ 0.56
TABLE XVIII: Sensitivity analysis of PPR topk.
PPR topk Cora PubMed ogbn-arxiv
64 69.54 $\pm$ 0.53 68.76 $\pm$ 1.31 69.67 $\pm$ 0.55
96 69.89 $\pm$ 0.48 69.05 $\pm$ 1.25 69.97 $\pm$ 0.46
128 70.11 $\pm$ 0.44 69.24 $\pm$ 1.21 70.22 $\pm$ 0.42
160 70.08 $\pm$ 0.47 69.17 $\pm$ 1.24 70.15 $\pm$ 0.45
192 69.92 $\pm$ 0.51 68.98 $\pm$ 1.28 70.01 $\pm$ 0.48
VI-D Ablation Studies
We conduct a series of ablation studies to verify the contributions of the designs in PGT under the linear probing setting.
Effect of pre-training objectives. Table XI shows the results, where “w/o feature recon.” represents that we only employ the local graph structure reconstruction objective in pre-training. Conversely, “w/o local graph recon.” signifies that only the feature reconstruction objective is used. It can be observed that for two downstream tasks, both proposed pre-training objectives contribute to performance. Comparatively, the removal of the feature reconstruction objective results in a larger performance degradation, implying the significance of node features for graph learning tasks.
Effect of key designs in inference. Table XII shows the influence of key proposed designs in inference. Our proposed method of employing PPR sampling for inference attains performance comparable to that achieved using all neighbor nodes. This substantiates the efficacy of PPR in selecting informative nodes without the necessity for a large receptive field. The notation “w/o proposed feature aug.” represents the absence of our proposed strategy of reusing the decoder for feature augmentation. Functioning as a plug-and-play component, it effectively improves performance across all datasets.
Efficiency analysis. Fig. 6 compares inference performance and time of different sampling methods, using the same number of nodes to be predicted (test set on each dataset) and the same hardware conditions. It can be observed that our method delivers the most favorable outcomes in both performance and speed. It is noteworthy that, in contrast to the efficiency of ClusterGCN sampling during the training phase, it does not exhibit advantages in inference. A possible explanation could be that the nodes requiring predictions constitute only a small fraction of all nodes. The GNN still needs to perform inference across all sub-graphs to obtain all prediction results. This further substantiates the significance of decoupling the sampling of labelled and auxiliary nodes.
VI-E Ablation Study on Feature Augmentation Strategies
To validate our proposed feature augmentation strategy, we compare it against several alternatives: no augmentation (original features only), feature dropout (10%), and adding Gaussian noise ($\mathcal{N}(0,0.1)$). As shown in Table XIII, our decoder-based augmentation consistently outperforms all other strategies. While common techniques like feature dropout and adding noise provide less benefit and can harm performance, our method yields consistent and meaningful improvements.
The effectiveness of our approach stems from its ability to leverage patterns learned during pre-training. Unlike random perturbations, the decoder generates features that are semantically consistent with the originals, acting as a form of ”learned feature denoising.” This process enhances feature representations by incorporating information from the graph’s global structure, confirming that the decoder captures meaningful patterns that benefit downstream tasks.
We further analyze how to best combine the original features ($X$) with the reconstructed features ($X^{\prime}$). We evaluate four strategies:
  • Original features only: Using only the original node features $X$.
  • Reconstructed features only: Using only the decoder-reconstructed features $X^{\prime}$.
  • Average pooling: Using $\frac{1}{2}(X+X^{\prime})$ (our proposed method).
  • Sum pooling: Using $(X+X^{\prime})$ without normalization.
Table XIV shows that combining features via average or sum pooling significantly outperforms using either original or reconstructed features alone. This confirms that the two feature sets provide complementary information, leading to more robust representations.
The performance of average and sum pooling is statistically similar. Theoretically, average pooling maintains the original feature scale, which can stabilize training. Sum pooling doubles the feature magnitude, which could require re-tuning other parameters. We adopt average pooling as our default strategy because it preserves the feature scale and performed slightly better in our experiments, though both are effective combination methods.
VI-F Validating The Choice of Transformers as The Backbone Architecture
To validate our choice of a transformer backbone, we compare it against a GNN-based approach across all experiments. As shown in Tables II, III, VI, VII, and X, the non-pretrained transformer consistently outperforms its GNN counterpart (No-pretrainGAT), even without pre-training. For instance, on the minors detection task (Table VI), the transformer baseline achieves ROC-AUC scores of 0.6771 and 0.8021 on Games A and B respectively, compared to 0.6621 and 0.7829 for the GAT baseline. This performance gap widens further after applying our pre-training framework.
We attribute this superiority to the transformer’s ability to: (1) handle heterophilic patterns common in industrial data; (2) remain robust to structural noise via self-attention; and (3) capture long-range dependencies without the over-smoothing issues that plague deep GNNs. Furthermore, as illustrated in Fig. 6, our transformer-based approach with PPR sampling is up to 12.9x faster in inference than full neighborhood aggregation methods while maintaining superior performance, addressing a critical requirement for industrial deployment.
VI-G Time Complexity Analysis
The complexity of our method comprises three main components: PPR computation, model training, and inference. For PPR computation, we use the efficient Forward Push algorithm with a complexity of approximately $O(k/\varepsilon)$ per seed node, which is far more scalable than the $O(|V|^{3})$ complexity of exact PPR. During training and inference, our model operates on fixed-length sequences. The training complexity for a batch of $b$ nodes is $O(b\times s^{2}\times d)$, where $s$ is the sequence length, and inference is $O(s^{2}\times d)$ per node. This is significantly more efficient than full-graph attention ($O(|V|^{2}\times d)$) or recursive neighborhood expansion.
Table XV details the practical computation times. For our 540M-node pre-training graph, computing PPR sequences for 54M seed nodes took approximately 8 hours on a 200-worker distributed cluster. On downstream tasks, this process took only 15-30 minutes. These measurements confirm the practicality of our approach for industrial applications requiring both training efficiency and low-latency inference.
VI-H Hyperparameter Sensitivity Analysis
To provide a more comprehensive understanding of our framework’s behavior, we conducted sensitivity analyses on three key hyperparameters: the mixing coefficient $\lambda$, mask rate, and PPR topk. We performed these experiments on three public datasets: Cora, PubMed, and ogbn-arxiv.
Sensitivity Analysis of $\lambda$. We varied $\lambda$ across the values 0.01,0.05,0.1,0.2,0.5{0.01,0.05,0.1,0.2,0.5} to understand its impact on model performance. The results are presented in Table XVI. These results demonstrate that while $\lambda=0.1$ yields optimal performance across all datasets, the model’s performance does not vary significantly across different values of $\lambda$. The maximum performance variation is approximately 0.4 percentage points across the tested range, indicating that our framework is not highly sensitive to this hyperparameter. This robustness to the $\lambda$ setting is advantageous from a practical perspective, as it suggests that precise tuning of this hyperparameter is not critical for achieving good performance with our framework across different datasets.
Sensitivity Analysis of Mask Rate. We varied the mask rate across the values 0.65, 0.75, 0.85, 0.90, 0.95 to evaluate its effect on performance. Table XVII presents the results. The results indicate that a mask rate of 0.85 yields the best performance across all datasets. This is consistent with findings in other masked modeling works that high mask rates (but not too extreme) often lead to optimal performance by forcing the model to learn more robust representations.
Sensitivity Analysis of PPR topk. We varied the PPR topk parameter across the values 64, 96, 128, 160, 192 to understand its impact on the model’s performance. Table XVIII shows the results. The results show that a PPR topk value of 128 consistently provides the best performance across all datasets. The model’s performance is relatively stable across different topk values, with only minor fluctuations.
VII conclusion